lilmami:

Determine the x- and y-intercepts of the linear equation. Type a numerical answer in the space provided. Type the x-intercept first, then the y-intercept second. Separate the intercepts with a comma. For example, if the x-intercept is (-10,0) and the y-intercept is (0,5), type -10,5. x - y = 3

lilmami:

help meh

snowflake0531:

okay, do you know what the slope-intercept form is? also, do you know what is the x andy intercept

lilmami:

no

snowflake0531:

Okay so the slope-intercept form is y=mx+b and the x-intercept is when y=0, or, when it intersects the x axis and the y-intercept is when x=0 or, when it intersects the y axis got that?

snowflake0531:

anyways, in x-y=3 first finding the x-intercept, plug in y as 0 so x-0=3, so x=3, so it would be (3,0) then, the y-intercept, plug in x as 0 0-y=3 y+3=0 y=-3 so, it would be (0,-3)

snowflake0531:

so, the x intercept: (3,0) y-intercept: (0,-3)

snowflake0531:

you could also use desmos, and make sure those are the points where it crosses the y-axis adn x-axis

snowflake0531:

do you get that? maybe? yes? no?

lilmami:

yes thx

snowflake0531:

yw

snowflake0531:

and thx for the medal

lilmami:

np